The length of a table is 5 meters greater than the width. The area of the table is 104m squared. What is the lenth and width?
Answer by stanbon(75887)

The length of a table is 5 meters greater than the width. The area of the table is 104m squared. What is the leng
th and width?
-----------------
let width be W
then length is W+5
----
Area = length * width
104 = w(w+5)
104 = w^2+5w
w^2+5w-104 = 0
Use Quadratic formula:
w = [-5+-sqrt(25-4*-104)]/2
Positive solution:
w = [-5+sqrt(441)]/2
etc.
---
Cheers,
Stan H.